approxmately 2.5kg of oranges are used to make 1liter of juice. the juice is poured into the 5 litres plastic bottles. determine the number of 5 l bottles of juice that can be made from 400kg of oranges


1
Expert's answer
2021-03-31T15:26:40-0400

Volume of juice: "V=\frac{400}{2.5}=160\;l."

Number of bottles: "N=\frac{V}{5}=\frac{160}{5}=32."